### v2

# Unveiling the inconsistency of the Proca theory with non-minimal coupling to gravity

Abstract
We study the degrees of freedom of the Proca theory, non-minimally coupled to gravity. In the Minkowski background, this theory propagates five degrees of freedom – a massive longitudinal mode, two massive vector ones, and two massless tensor modes. At first sight, the non-linear coupling between the metric perturbations and the vector field indicates that both longitudinal and tensor modes become strongly coupled, at the same scale. This would imply that no matter how small the photon mass is if non-minimal coupling is taken into account, gravitational waves would necessarily be strongly coupled. We show that the way out of this inconsistency is through the introduction of the disformal coupling to the metric perturbations that resemble the vector-type disformal transformations. This way, the unphysical coupling between the two types of modes can be avoided, rendering the model consistent. As a result, we show that only the longitudinal modes enter a strong coupling regime, while both tensor and transverse modes remain weakly coupled at all scales up to the Planck length. Finally, we show how this cures recently reported runaway modes in curved spacetime.
1 Introduction
One of the most important contemporary puzzles underlying the structure of our Universe is – What is dark matter? The existence of this component, which should account for the missing mass of the Universe is strongly supported by the observational data. Yet, the theory that explains its origin remains unknown.
Among numerous proposals for models of dark matter, the theory of the dark photon is particularly interesting. This theory assumes that the photon is described by a massive and stable vector field, and can be detectable due to the kinetic mixing with the ordinary photon [1, 2]. Because of this mixing, these dark matter particles can be produced in the early Universe, for a sufficiently large mass. Gravitational particle production – the mechanism in which the particles are produced due to the curvature of the space-time – can account for their production in the early Universe if, in contrast, the photon is ultra-light [3]. To date, this mechanism has brought to us the understanding of the origin of galaxies and the large-scale structure, tracing its seeds back to the quantum fluctuations during an inflationary era [4, 5, 6, 7, 8, 9, 10], and is even responsible for the Hawking radiation of black holes [11]. Notably, with it, the dark photons can be produced independently of any coupling to the standard-model particles, making the particles completely dark [12].
The theory underlying the physics of these dark matter particles is known as the Proca theory [13]. When considered in curved space-time, one way to study it is by assuming the minimal coupling between the vector field and the curvature, through a simple generalization of the volume element in the action and promotion of the partial derivatives to the covariant ones. The cosmological implications of this theory were widely studied in [14, 15, 16]. Recently, the possibility of having non-minimal couplings between a vector field and the curvature tensor has become very attractive as well. This natural extension of the model, whose quantization was discussed in [17, 18], was particularly intriguing due to its possibility to realize inflation due to a vector field and generate the primordial magnetic fields [19, 20, 21, 22, 23]. Moreover, it can influence the gravitational particle production of the dark photon [24, 12, 25, 26]. In this work, we will also consider Proca theory, non-minimally coupled to gravity, studying it from a complementary perspective that arises from the recent developments in massive gauge theories.
Intuitively, we can think of the Proca theory as a small modification of Maxwell’s electrodynamics, that amounts to adding a mass term by hand. Then, per the principle of continuity, one might expect that in the limit when this mass is set to zero, all observables of the modified theory should smoothly go back to the original ones [27]. In general, however, it took several decades to show that this principle holds in massive gauge theories. The two particularly interesting examples are massive Yang-Mills theory (mYM) with mass added by hand, and massive linearized gravity [28, 29]. Studying the two with conventional methods could lead to the conclusion that they might not be realized in nature. In the case of mYM theory, when calculating the corrections to the propagator of the non-Abelian vector field it was shown that the perturbative series are singular in mass [30, 31, 32, 33]. For the mGR case, the situation was even more drastic. It yielded predictions that were different from those of Einstein’s relativity and remained such in the massless limit. As pointed out in [30, 34], the occurrence of this perturbative discontinuity in mass was due to the appearance of the extra degrees of freedom (dof), the longitudinal modes, that failed to decouple from the remaining dof in the massless limit. However, as pointed out in [35] in the context of massive gravity, once the non-linear terms become of the same order as the linear ones, the perturbation theory for the longitudinal mode breaks down. Subsequently, on smaller distances than this Vainshtein radius, the longitudinal modes are strongly coupled and one obtains predictions of linearized gravity up to small corrections [36, 37]. The possibility of the smooth massless limit for the mYM was conjectured in [38] as well, and later shown in [39]. The Vainsthein mechanism – mechanism which is responsible for this behavior of the modes absent in massless theories was also generalized to other field theories [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. Additionally, it also holds for the Proca theory, in the presence of the quartic and cubic self-interactions, and generalized Proca theories [52, 53, 50, 54].
The main question that we aim to answer in this paper is – Does Proca theory non-minimally coupled to gravity have strong coupling, and if yes, how strongly coupled the theory is? Based on the results in massive gauge theories, there are three probable outcomes: First, the theory could be already well-behaved in the massless limit, despite the presence of the longitudinal modes. In contrast, the second two possibilities contain perturbative series that are singular in mass, and consequently, which have a strong coupling scale. However, either only the longitudinal mode becomes strongly coupled, or the remaining modes of the theory become strongly coupled as well. From the perspective of self-interacting theory, the first case is natural. In cubic or quartic self-interactions, the transverse modes, dof which are also present in Maxwell’s electrodynamics, remain weakly coupled even after the perturbation theory for the longitudinal modes breaks down. The second outcome, in which the transverse modes, or tensor modes – modes of the metric perturbations – become strongly coupled, however, would be unphysical. It would imply that in the presence of non-minimal coupling, one could either not be able to recover electromagnetism or discover gravitational waves.
By studying the Proca theory, non-minimally coupled to gravity in a flat background, and taking into account the metric perturbations around it, we will show that the last case takes place – both tensor and longitudinal modes become strongly coupled, yielding a physically inconsistent result. Moreover, we will show that the way to resolve this ambiguity lies in the introduction of the disformal coupling to the metric perturbations simultaneously. Such disformal transformations, which modify the original metric by adding a bilinear term in the external fields or their field-strength have gained a lot of attention in cosmology, and construction of new theories of gravity [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. In addition, we will show that the production of runaway modes – modes with negative frequency in curved spacetime – recently reported in [26], can be cured with the same transformations.
The paper is organized in the following way. First, we will review the free Proca theory and linearized gravity in a flat space-time background, and identify potential sources of the strong coupling. Then, we will take into account the non-linear terms, and analyze each type of the non-minimal coupling separately. This will be followed by the resolution of the problem that appears for the tensor modes, and subsequently the resolution of the runaway in the curved space-time. Finally, we will conclude with a discussion of the results and future prospects of the theory.
2 The basics of linearized gravity and Proca theory
The Proca theory, the theory of a massive vector field, which is non-minimally coupled to gravity, is described by the following action [^1] We use the metric signature (-,+,+,+).:
$S=S_{EH}+S_{P}+S_{nmin},$ (1)
where
$\begin{split}&S_{EH}=\frac{M_{pl}^{2}}{2}\int d^{4}x\sqrt{-g}R\\ &S_{P}=\int d^{4}x\sqrt{-g}\left(-\frac{1}{4}g^{\mu\alpha}g^{\nu\beta}F_{\mu\nu}F_{\alpha\beta}-\frac{m^{2}}{2}g^{\mu\nu}A_{\mu}A_{\nu}\right)\\ &S_{nmin}=-\frac{1}{2}\int d^{4}x\sqrt{-g}\left(\alpha Rg^{\mu\nu}A_{\mu}A_{\nu}+\beta R^{\mu\nu}A_{\mu}A_{\nu}\right).\end{split}$
The goal of this section is to analyze the dof of this theory, starting with the flat background. For this, let us perturb the metric in the following way:
$g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu},$ (2)
where $\eta_{\mu\nu}$ is the Minkowski metric. Then, we find that the Lagrangian density corresponding to (1) at linearized level is given by:
$\mathcal{L}=\mathcal{L}_{EH}+\mathcal{L}_{P},$ (3)
where
$\begin{split}\mathcal{L}_{EH}=&\frac{M_{pl}^{2}}{4}\left(-\partial_{\alpha}h^{\mu\nu}\partial^{\alpha}h_{\mu\nu}+2\partial_{\mu}h^{\mu\nu}\partial_{\alpha}h_{\nu}^{\alpha}-2\partial_{\nu}h\partial_{\mu}h^{\mu\nu}+\partial_{\mu}h\partial^{\mu}h\right),\\ \mathcal{L}_{P}=&-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{m^{2}}{2}A_{\mu}A^{\mu}.\end{split}$ (4)
Here, the indices are raised and lowered the indices with the Minkowski metric.
Let us now first study the dof in gravity. For this, we will decompose the metric perturbations according to the spatial rotations:
$\begin{split}&h_{00}=2\phi\\ &h_{0i}=B_{,i}+S_{i},\qquad\qquad S_{i,i}=0\\ &h_{ij}=2\psi\delta_{ij}+2E_{,ij}+F_{i,j}+F_{j,i}+h_{ij}^{T},\qquad\qquad F_{i,i}=0,\quad h_{ij,i}^{T}=0,\quad h_{ii}^{T}=0,\end{split}$ (5)
where $,i=\frac{\partial}{\partial x^{i}}$. Then, the gravitational contribution of (4) becomes:
$\begin{split}\mathcal{L}_{EH}=&M_{pl}^{2}\left(2\phi\Delta\psi-3\dot{\psi}\dot{\psi}-\psi\Delta\psi-2\psi\Delta\dot{\sigma}\right)-\frac{M_{pl}^{2}}{4}V_{i}\Delta V_{i}-\frac{M_{pl}^{2}}{8}\partial_{\mu}h_{ij}^{T}\partial^{\mu}h_{ij}^{T},\end{split}$ (6)
where
$\sigma=B-\dot{E},\qquad\text{and}\qquad V_{i}=S_{i}-\dot{F}_{i}.$ (7)
In the above relations, the dot denotes a derivative with respect to the time. We can notice that $\phi$ and $S_{i}$ are not propagating. By varying the action with respect to them, we find the following equations respectively:
$\Delta\psi=0\qquad\text{and}\qquad\Delta V_{i}=0,$ (8)
whose solutions are given by:
$\psi=0\qquad\text{and}\qquad V_{i}=0.$ (9)
Thus, at the linearized level, we confirm that the gravitational contribution consists only of the two tensor modes, which are given by:
$\mathcal{L}_{EH}=-\frac{M_{pl}^{2}}{8}\partial_{\mu}h_{ij}^{T}\partial^{\mu}h_{ij}^{T}.$ (10)
Let us now look at the vector part. Similarly to the metric perturbations, we will separate the temporal and spatial components of the vector field, and further decompose the spatial part according to the following:
$A_{i}=A_{i}^{T}+\chi_{,i},\qquad A_{i,i}^{T}=0.$ (11)
Then, the Lagrangian density corresponding to the vector field becomes:
$\begin{split}\mathcal{L}_{P}=\frac{1}{2}&\left[A_{0}\left(-\Delta+m^{2}\right)A_{0}+2A_{0}\Delta\dot{\chi}-\left(\dot{\chi}\Delta\dot{\chi}-m^{2}\chi\Delta\chi\right)\right.\\ &\left.+\left(\dot{A_{i}^{T}}\dot{A_{i}^{T}}-A_{i,j}^{T}A_{i,j}^{T}-m^{2}A_{i}^{T}A_{i}^{T}\right)\right]\end{split}$ (12)
We can see that the temporal component is not-propagating, and satisfies the constraint:
$(-\Delta+m^{2})A^{0}=-\Delta\dot{\chi},$ (13)
Substituting its solution
$A_{0}=\frac{-\Delta}{-\Delta+m^{2}}\dot{\chi}.$ (14)
back to the action, we find:
$\mathcal{L}_{P}=-\frac{1}{2}\left[A_{i}^{T}(-\Box+m^{2})A_{i}^{T}+\chi(-\Box+m^{2})\frac{m^{2}(-\Delta)}{-\Delta+m^{2}}\chi\right].$ (15)
Thus, we have found two transverse vector modes, $A_{i}^{T}$, and a longitudinal mode.
At the linearized level, the dof of the de Broglie-Proca theory are decoupled. However, once one considers higher-order contributions, that arise due to the coupling of the Proca theory to gravity, as well as the non-minimal coupling, the modes can influence each other’s behavior. In this work, our aim is to investigate these corrections, and ask – At which scale can the perturbation theory break down?
The main tool that will help us to answer this question is the minimal amplitude of quantum fluctuations – equal time, two-point colleration function, which is the direct consequence of Heisenberg’s uncertainty principle. For a normalized quantum field, the typical amplitude of the quantum fluctuations for scales $k$ corresponding to length scales $L\sim\frac{1}{k}$ is of the order of $\left(\frac{k^{3}}{\omega_{k}}\right)^{\frac{1}{2}}$ [75].
By canonically normalizing the tensor modes according to:
$h_{(n)ij}^{T}=\frac{M_{pl}}{2}h_{ij}^{T},$ (16)
we thus find that quantum fluctuations of the original modes behave as:
$\delta h^{T}_{L}\sim\frac{1}{M_{pl}L}.$ (17)
Similarly, the longitudinal modes need to be canonically normalized as well, according to:
$\chi_{n}=\sqrt{\frac{-\Delta m^{2}}{-\Delta+m^{2}}}\chi,$ (18)
while the transverse modes are already in the canonical form. For scales $k^{2}\gg m^{2}$, we thus find that the minimal amplitude of quantum fluctuations for these modes is respectively given by:
$\delta\chi_{L}\sim\frac{1}{mL}\qquad\text{and}\qquad\delta A_{L}^{T}\sim\frac{1}{L}.$ (19)
In contrast to the transverse modes, we can notice that the quantum fluctuations of the longitudinal ones are singular in mass. In the presence of the self-interactions of the Proca field, it is well known that these introduce a breakdown in the perturbation theory [50]. Let us now investigate if the same will hold true in the presence of the non-minimal coupling.
3 Investigating the Strong Coupling
In the previous section, we have seen that each type of mode behaves differently. In contrast to the transverse modes, which are only dependent on the length scale, the amplitude of the quantum fluctuations for the tensor modes is inversely proportional to the Planck mass. Most importantly, the amplitude of quantum fluctuations for the longitudinal modes is singular in the mass of the vector field. Due to this behavior, in the presence of self-interactions such as $\left(A_{\mu}A^{\mu}\right)^{2}$, the longitudinal modes will necessarily enter a strong coupling regime [50] (see also [39] for the non-Abelian generalization).
In this section, we will investigate if the longitudinal modes of the non-minimally coupled theory, described in (1) become strongly coupled as well. For this, let us go beyond the linear analysis. By expanding the metric perturbations around the flat background as in the previous chapter, we find the following Lagrangian density:
$\mathcal{L}=\mathcal{L}_{EH}+\mathcal{L}_{P}+\mathcal{L}_{\alpha int}+\mathcal{L}_{\beta int},$ (20)
where
$\begin{split}\mathcal{L}_{EH}=&\frac{M_{pl}^{2}}{4}\left(2h^{\mu\nu}_{,\mu}h_{\alpha\nu}^{,\alpha}-h_{\mu\nu,\alpha}h^{\mu\nu,\alpha}-2h^{\mu\nu}_{,\mu}h_{,\nu}+h_{,\mu}h^{,\mu}\right)\\ &+\frac{M_{pl}^{2}}{2}\left(-\frac{1}{4}h^{\alpha\beta}h_{\mu\nu}h^{\mu\nu}_{,\alpha\beta}-\frac{1}{8}h_{\mu\nu}h^{\mu\nu}h^{\alpha\beta}_{,\alpha\beta}-\frac{1}{4}h^{\alpha\beta}h_{\beta\nu}h_{\alpha,\delta}^{\nu,\delta}-h^{\alpha\beta}h_{\alpha,\mu}^{\mu}h^{\nu}_{\beta,\nu}\right.\\ &\left.+\frac{1}{2}h^{\alpha\beta}h_{\beta\nu,\mu}h^{\mu,\nu}_{\alpha}-h^{\alpha\beta}_{,\mu}h_{\alpha}^{\mu}h^{\nu}_{\beta,\nu}+\frac{1}{8}hh_{\mu\nu}h^{\mu\nu,\gamma}_{,\gamma}+\frac{3}{16}h_{\mu\nu}h^{\mu\nu}h_{,\gamma}^{,\gamma}+h^{\alpha\beta}h_{\alpha,\mu}^{\mu}h_{,\beta}\right.\\ &\left.+\frac{1}{2}hh^{\mu\nu}_{,\mu}h^{\alpha}_{\nu,\alpha}-\frac{1}{4}hh_{\alpha\nu,\mu}h^{\alpha\mu,\nu}+\frac{1}{4}hh^{\alpha\beta}h_{,\alpha\beta}-\frac{1}{16}h^{2}h^{,\mu}_{,\mu}\right),\\ \mathcal{L}_{P}=&-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{m^{2}}{2}A_{\mu}A^{\mu}\\ &+\frac{1}{2}h^{\mu\alpha}F_{\mu\nu}F_{\alpha}^{\;\nu}+\frac{m^{2}}{2}h^{\mu\nu}A_{\mu}A_{\nu}+\frac{1}{2}h\left(-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{m^{2}}{2}A_{\mu}A^{\mu}\right),\\ \mathcal{L}_{\alpha int}=&-\frac{\alpha}{2}\left(h^{\mu\nu}_{,\mu\nu}-\Box h\right)A_{\mu}A^{\mu},\\ \mathcal{L}_{\beta int}=&-\frac{\beta}{4}A_{\mu}A_{\nu}\left(2h^{\nu\alpha,\mu}_{,\alpha}-h^{,\mu\nu}-\Box h^{\mu\nu}\right)\end{split}$
In the following analysis, we will decompose the metric and the vector field according to the spatial rotations according to (5) and 11) respectively, and apply the conformal gauge $E=B=0$, together with $F_{i}=0$.
In order to gain insights into the structure of the interactions, we will study each non-minimal coupling case separately.
3.1 $\beta$ coupling case
As a first step, let us set $\alpha=0$, and consider only the coupling of the vector fields with the Ricci tensor. In this case, the most important terms in the Lagrangian density are given by:
$\mathcal{L}=\mathcal{L}_{EH}+\mathcal{L}_{P}+\mathcal{L}_{\beta int},$ (21)
where
$\begin{split}\mathcal{L}_{EH}\sim&M_{pl}^{2}\left(2\phi\Delta\psi-3\dot{\psi}\dot{\psi}-\psi\Delta\psi\right)-\frac{M_{pl}^{2}}{4}S_{i}\Delta S_{i}-\frac{M_{pl}^{2}}{8}\partial_{\mu}h_{ij}^{T}\partial^{\mu}h_{ij}^{T}\\ &-\frac{M_{pl}^{2}}{4}\left(\frac{1}{2}h_{ij}^{T}h_{ik}^{T}h_{jk,\mu}^{T,\mu}+\frac{1}{2}h_{ij}^{T}h_{kl}^{T}h_{kl,ij}^{T}-h_{ij}^{T}h_{il,k}^{T}h_{jk,l}^{T}\right),\\ \\ \mathcal{L}_{P}=&\frac{1}{2}\left[A_{0}\left(-\Delta+m^{2}\right)A_{0}+2A_{0}\Delta\dot{\chi}-\left(\dot{\chi}\Delta\dot{\chi}-m^{2}\chi\Delta\chi\right)\right.\\ &\left.+\left(\dot{A_{i}^{T}}\dot{A_{i}^{T}}-A_{i,j}^{T}A_{i,j}^{T}-m^{2}A_{i}^{T}A_{i}^{T}\right)\right]\\ &+\psi\left[\left(A_{j,i}^{T}-A_{i,j}^{T}\right)\left(A_{j,i}^{T}-A_{i,j}^{T}\right)-\left(\dot{A}_{i}-A_{0,i}\right)\left(\dot{A}_{i}-A_{0,i}\right)+m^{2}A_{i}A_{i}\right]\\ &+\frac{1}{2}(3\psi-\phi)\left[\dot{A}_{i}\dot{A}_{i}-A_{j,i}A_{j,i}+A_{j,i}A_{i,j}+A_{0,i}A_{0,i}-2A_{0,i}\dot{A}_{i}+m^{2}A_{0}^{2}-m^{2}A_{i}A_{i}\right]\\ &+\phi\left[\left(\dot{A}_{i}-A_{0,i}\right)\left(\dot{A}_{i}-A_{0,i}\right)+m^{2}A_{0}^{2}\right]-S_{i}\left[\left(\dot{A}_{j}-A_{0,j}\right)\left(A^{T}_{j,i}-A^{T}_{i,j}\right)+m^{2}A_{0}A_{i}\right]\\ &+\frac{1}{2}h_{ij}^{T}\left[\left(A_{k,i}^{T}-A_{i,k}^{T}\right)\left(A_{k,j}^{T}-A_{j,k}^{T}\right)-\left(A_{0,i}-\dot{A}_{i}\right)\left(A_{0,j}-\dot{A}_{j}\right)+m^{2}A_{i}A_{j}\right]\\ \\ \mathcal{L}_{\beta int}=&\frac{\beta}{2}\left[A_{0}^{2}\left(\Delta\phi+3\ddot{\psi}\right)-A_{0}A_{i}\left(\Delta S_{i}+4\dot{\psi}_{,i}\right)\right]\\ &+\frac{\beta}{4}A_{i}A_{j}\left(2\psi_{,ij}+2\Box\psi\delta_{ij}-2\phi_{,ij}+\dot{S}_{j,i}+\dot{S}_{i,j}+\Box h_{ij}^{T}\right)\end{split}$
Here, we have kept $A_{i}$ in some parts in order to keep expressions simpler.
One can notice that we have neglected the cubic terms of $\psi$ ad $S_{i}$, as well as their combinations that involve $\phi$ in $\mathcal{L}_{EH}$. These fields would yield at least quartic terms that mix with the tensor modes. The reason for this lies in the constraints. By varying with respect to $\phi$, one finds that $\psi$ satisfies to the leading order:
$\begin{split}\psi=&\frac{\beta}{4M_{pl}^{2}}\left[\frac{\partial_{i}\partial_{j}}{\Delta}\left(A_{i}A_{j}\right)-A_{0}A_{0}\right]-\frac{1}{2M_{pl}^{2}}\frac{1}{\Delta}\left[\left(\dot{A}_{i}-A_{0,i}\right)\left(\dot{A}_{i}-A_{0,i}\right)+m^{2}A_{0}^{2}\right]\\ &+\frac{1}{4M_{pl}^{2}}\frac{1}{\Delta}\left[\dot{A}_{i}\dot{A}_{i}-A_{j,i}A_{j,i}+A_{j,i}A_{i,j}+A_{0,i}A_{0,i}-2A_{0,i}\dot{A}_{i}+m^{2}A_{0}^{2}-m^{2}A_{i}A_{i}\right]\end{split}$ (22)
The vector mode $S_{i}$ is constrained as well. The leading order solution for its constraint is given by:
$\begin{split}S_{k}=&\frac{\beta}{M_{pl}^{2}}P_{ki}^{T}\left[-A_{0}A_{i}+\frac{1}{2}\frac{1}{\Delta}\partial_{j}\partial_{0}\left(A_{j}A_{j}\right)\right]-\frac{2}{M_{pl}^{2}}P_{ki}^{T}\left[\left(\dot{A}_{j}-A_{0,j}\right)\left(A^{T}_{j,i}-A^{T}_{i,j}\right)-m^{2}A_{0}A_{i}\right],\end{split}$ (23)
where
$P_{ij}^{T}=\delta_{ij}-\frac{\partial_{i}\partial_{j}}{\Delta}$ (24)
is the transverse projector.
By substituting these expressions back to (21), we find up to cubic order:
$\mathcal{L}=\mathcal{L}_{EH}+\mathcal{L}_{P}+\mathcal{L}_{\beta int},$ (25)
where
$\begin{split}\mathcal{L}_{EH}=&-\frac{M_{pl}^{2}}{8}\partial_{\mu}h_{ij}^{T}\partial^{\mu}h_{ij}^{T}-\frac{M_{pl}^{2}}{4}\left(\frac{1}{2}h_{ij}^{T}h_{ik}^{T}h_{jk,\mu}^{T,\mu}+\frac{1}{2}h_{ij}^{T}h_{kl}^{T}h_{kl,ij}^{T}-h_{ij}^{T}h_{il,k}^{T}h_{jk,l}^{T}\right),\\ \\ \mathcal{L}_{P}=&\frac{1}{2}\left[A_{0}\left(-\Delta+m^{2}\right)A_{0}+2A_{0}\Delta\dot{\chi}-\left(\dot{\chi}\Delta\dot{\chi}-m^{2}\chi\Delta\chi\right)\right.\\ &\left.+\left(\dot{A_{i}^{T}}\dot{A_{i}^{T}}-A_{i,j}^{T}A_{i,j}^{T}-m^{2}A_{i}^{T}A_{i}^{T}\right)\right]\\ &+\frac{1}{2}h_{ij}^{T}\left[\left(A_{k,i}^{T}-A_{i,k}^{T}\right)\left(A_{k,j}^{T}-A_{j,k}^{T}\right)-\left(A_{0,i}-\dot{A}_{i}\right)\left(A_{0,j}-\dot{A}_{j}\right)+m^{2}A_{i}A_{j}\right]\\ \\ \mathcal{L}_{\beta int}=&\frac{\beta}{4}A_{i}A_{j}\Box h_{ij}^{T}\end{split}$
We can notice that similarly to the free case, the temporal component of the vector field is not propagating. By varying the action with respect to it, we find:
$(-\Delta+m^{2})A_{0}=-\Delta\dot{\chi}-h_{ij}^{T}\left(A_{0,ij}-\dot{A}_{j,i}\right)$ (26)
By solving it perturbatively, and substituting the solution back to the action, we find:
$\begin{split}\mathcal{L}\sim&-\frac{M_{pl}^{2}}{8}\partial_{\mu}h_{ij}^{T}\partial^{\mu}h^{T}_{ij}-\frac{1}{2}A_{i}^{T}(-\Box+m^{2})A_{i}^{T}-\frac{1}{2}\chi(-\Box+m^{2})\frac{m^{2}(-\Delta)}{-\Delta+m^{2}}\chi\\ &+\frac{\beta}{4}\left(A_{i}^{T}+\chi_{,i}\right)\left(A_{j}^{T}+\chi_{,j}\right)\Box h_{ij}^{T}+\frac{1}{2}h_{ij}^{T}\left[\left(A_{k,j}^{T}-A_{j,k}^{T}\right)\left(A_{k,i}^{T}-A_{i,k}^{T}\right)-\dot{A}_{i}^{T}\dot{A}_{j}^{T}\right]\\ &\frac{m^{2}}{2}h_{ij}^{T}\left(\chi_{,i}\chi_{,j}+2\chi_{,i}A_{j}^{T}\right),\end{split}$ (27)
where have kept the most relevant terms.
Let us now study the equations of motion of the propagating modes. The most important contributions to the equation of motion of the longitudinal modes are given by:
$(-\Box+m^{2})m^{2}\chi\sim-\frac{\beta}{2}\partial_{i}\left[\chi_{,j}\Box h_{ij}^{T}\right]-m^{2}\partial_{i}\left(h_{ij}^{T}\chi_{,j}\right)$ (28)
For the transverse modes, we find:
$\begin{split}(-\Box+m^{2})A_{l}^{T}=P_{li}^{T}&\left[\frac{\beta}{2}\left(A_{j}^{T}+\chi_{,j}\right)\Box h_{ij}^{T}+\dot{h}_{ij}^{T}\dot{A}_{j}^{T}\right.\\ &\left.-h_{ij}^{T}\left(A_{i,jk}^{T}-A_{j,ik}^{T}\right)+\partial_{k}\left(h_{ij}^{T}\left(A_{k,j}^{T}-A_{j,k}^{T}\right)\right)\right]\end{split}$ (29)
Finally, for the transverse modes, the leading order contributions are given by:
$\Box h_{ij}^{T}\sim-\frac{\beta}{M_{pl}^{2}}P_{ijkl}^{T}\Box\left(\chi_{,k}\chi_{,l}\right),$ (30)
where
$\begin{split}P_{mnij}^{T}&=\frac{1}{2}\left(\delta_{im}\delta_{jn}+\delta_{jm}\delta_{in}\right)-\frac{1}{2}\delta_{ij}\left(\delta_{mn}-\frac{\partial_{k}\partial_{l}}{\Delta}\delta_{km}\delta_{ln}\right)\\ &+\frac{1}{\Delta}\left[\frac{1}{2}\delta_{mn}\partial_{i}\partial_{j}+\frac{1}{2}\delta_{mk}\delta_{ln}\frac{\partial_{i}\partial_{j}\partial_{k}\partial_{l}}{\Delta}-\delta_{im}\delta_{ln}\partial_{l}\partial_{j}-\delta_{jm}\delta_{ln}\partial_{l}\partial_{i}\right]\end{split}$ (31)
is the transverse-traceless projector. Let us evaluate the above equations perturbatively:
$\chi=\chi^{(0)}+\chi^{(1)}+...\qquad A_{i}^{T}=A_{i}^{T(0)}+A_{i}^{T(1)}+...\qquad h_{ij}^{T}=h_{ij}^{T(0)}+h_{ij}^{T(1)}+...,$ (32)
with
$(-\Box+m^{2})\chi^{(0)}=0\qquad(-\Box+m^{2})A_{i}^{T(0)}=0\qquad\Box h_{ij}^{T(0)}=0.$ (33)
In this case, the first-order corrections for the longitudinal modes are given by:
$(-\Box+m^{2})m^{2}\chi^{(1)}=-m^{2}\partial_{i}\left(h_{ij}^{T(0)}\chi_{,j}^{(0)}\right)$ (34)
By taking into account (19) and (17), for scales $k^{2}\sim\frac{1}{L^{2}}\gg m^{2}$, we can evaluate them as
$\chi^{(1)}\sim\frac{1}{mM_{pl}L^{2}}.$ (35)
They become of the same order as the linear terms $\chi^{(0)}$ at the Planck scale:
$L_{\chi}\sim\frac{1}{M_{pl}}.$ (36)
Similarly, the first-order corrections to the transverse modes can be evaluated as:
$A_{i}^{T(1)}\sim\frac{1}{M_{pl}L^{2}},$ (37)
and become of the same order as the linear term at the same scale as the longitudinal mode.
The tensor modes, however, indicate strong coupling, that appears sooner than the Planck mass is reached. Their first-order corrections can be estimated as:
$h_{ij}^{(1)}\sim\frac{\beta}{M^{2}_{pl}m^{2}L^{4}}.$ (38)
Therefore, they become of the same order as the linear term $h_{ij}^{T(0)}\sim\frac{1}{M_{pl}L}$ at the following scale:
$L_{\beta str}\sim\left(\frac{\beta}{M_{pl}m^{2}}\right)^{1/3}.$ (39)
Once For the longitudinal modes, this means that the first term on the left-hand side of (28) is no longer vanishing. Once the scale $L_{\beta str}$ is reached, their first-order corrections should be evaluated as:
$\chi^{(1)}\sim\frac{\beta}{m^{3}M_{pl}L^{4}}.$ (40)
Due to them, the longitudinal mode becomes strongly coupled as well, as soon as one reaches the scale $L_{\beta str}$. Moreover, by studying the quartic order contributions, one can find that terms of the order $\mathcal{O}\left(\frac{\beta^{2}}{M_{pl}^{2}L^{6}}\chi^{4}\right)$ appear in the Lagrangian density, which is of the same order as the cubic contributions when evaluated at the strong coupling scale.
Therefore, in this section, we have found that both tensor modes and longitudinal modes become strongly coupled once the scale reaches $L_{\beta str}$. However, this is an unphysical result. It would imply that no matter how small the mass of the photon is, gravitational waves would be strongly coupled if non-minimal coupling is taken into account.
In the following section, we will first study the other non-minimal term, and then show how to avoid this inconsistency.
3.2 $\alpha$ coupling case
In the previous section we have seen that due to the $\beta$ interaction, both tensor and the longitudinal modes become strongly coupled. Let us now study the coupling that involves only the Ricci scalar. First of all, we can notice that in this case, at cubic order, only the gravitational potentials are included in the interaction:
$\mathcal{L}_{\alpha int}=-\alpha\left(\Delta\phi+3\ddot{\psi}-2\Delta\psi\right)A_{\mu}A^{\mu}$ (41)
Thus, in this order, we can safely avoid the danger of tensor modes becoming strongly coupled, in contrast to the previous case. Therefore, the only question that remains to be answered is if the longitudinal modes will enter a strong coupling regime at quartic order. For simplicity, let us set the tensor and transverse vector modes to zero. Then, the most dominant terms in the Lagrangian are given by:
$\mathcal{L}=\mathcal{L}_{EH}+\mathcal{L}_{P}+\mathcal{L}_{\alpha int},$ (42)
where
$\begin{split}\mathcal{L}_{EH}=&M_{pl}^{2}\left(2\phi\Delta\psi-3\dot{\psi}\dot{\psi}-\psi\Delta\psi\right)\\ \mathcal{L}_{P}=&\frac{1}{2}\left[A_{0}\left(-\Delta+m^{2}\right)A_{0}+2A_{0}\Delta\dot{\chi}-\left(\dot{\chi}\Delta\dot{\chi}-m^{2}\chi\Delta\chi\right)\right]\\ \mathcal{L}_{\alpha int}=&-\alpha\left(\Delta\phi+3\ddot{\psi}-2\Delta\psi\right)\left(\chi_{,i}\chi_{,i}-A_{0}A_{0}\right)\end{split}$ (43)
One should note that we have neglected also higher orders of the gravitational potentials as they would generate terms that are in higher order than quartic ones. Moreover, we have neglected the interactions that arise from the $\mathcal{L}_{P}$, as these will be multiplied by the mass term after resolving the constraints, and thus be subdominant.
Let us now study the equations of motion. By varying with respect to $\phi$, we find:
$2M_{pl}^{2}\Delta\psi=\alpha\Delta\left(\chi_{,i}\chi_{,i}-A_{0}A_{0}\right),$ (44)
which can be solved as:
$\psi=\frac{\alpha}{2M_{pl^{2}}}\left(\chi_{,i}\chi_{,i}-A_{0}A_{0}\right).$ (45)
By varying with respect to $\psi$, we further find:
$2M_{pl}\Delta\phi+6M_{pl}^{2}\ddot{\psi}-2M_{pl}^{2}\Delta\psi-3\alpha\partial_{0}^{2}\left(\chi_{,i}\chi_{,i}-A_{0}A_{0}\right)+2\alpha\Delta\left(\chi_{,i}\chi_{,i}-A_{0}A_{0}\right)=0$ (46)
By substituting (45) into the above equation, we find:
$\phi=-\psi$ (47)
within our approximation. Let us now consider the $A_{0}$ constraint. By varying the action with respect to this component, we find:
$(-\Delta+m^{2})A_{0}-6\alpha\Box\psi=-\Delta\dot{\chi},$ (48)
where we have used (47). In the limit $k^{2}\gg m^{2}$, the solution of the above constraint is given by:
$A_{0}\sim\frac{-\Delta}{-\Delta+m^{2}}\dot{\chi}+\frac{3\alpha^{2}}{M_{pl}^{2}}\frac{1}{-\Delta+m^{2}}\left[\dot{\chi}\Box(\chi_{,i}\chi_{,i}-\dot{\chi}\dot{\chi})\right]$ (49)
Finally, the longitudinal mode satisfies the following equation of motion:
$\Delta\ddot{\chi}+m^{2}\Delta\chi-\Delta\dot{A}_{0}-6\alpha\partial_{i}\left(\chi_{,i}\Box\psi\right)=0,$ (50)
where we have used (47) again. Substituting in addition the solution for the constraint of the temporal component, this equation can be written on scales $k^{2}\gg m^{2}$ as:
$-m^{2}(-\Box+m^{2})\chi\sim\frac{3\alpha^{2}}{M_{pl}^{2}}\partial_{\mu}\left[\chi^{,\mu}\Box(\chi_{,\alpha}\chi^{,\alpha})\right].$ (51)
Clearly, the term on the right-hand side is a non-vanishing one even within the perturbation theory. Thus, it yields a strong coupling for the longitudinal mode, with the scale given by:
$L_{\alpha}\sim\left(\frac{\alpha}{M_{pl}m^{2}}\right)^{1/3}$ (52)
We can notice that this scale arrives in the same manner as the strong coupling scale of the $\beta$ case, with only the coupling interchanged. However, while for the $\beta$ case we found that both tensor and longitudinal modes become strongly coupled, here only the longitudinal modes enter the strong coupling regime. This is a standard property of Proca theory and is thus not as problematic. Moreover, due to the scaling of the remaining modes – the vector and tensor ones – we can be sure that this will be the most dominant term.
4 How to cure the tensor modes?
In the previous sections, we have studied the $\alpha$ and $\beta$ models, which involved the coupling of the vector field with the Ricci scalar and the Ricci tensor respectively. In the case of the $\alpha$ coupling, we have found that only the longitudinal modes become strongly coupled, due to the quartic self-interaction that appears after one solves the constraints. In contrast, the coupling with the Ricci tensor proved to be troubling. There, we have found that both tensor and longitudinal modes become strongly coupled, which is a physically unacceptable result.
Therefore, one might accept that this is the case, and exclude such interaction in the further models. In this work, we will, however, try to resolve this puzzle in another way. Notably, a similar problem can be encountered in massive Yang-Mills theory, with mass added by hand [39]. If one decomposes the spatial part of this vector field only linearly, in a way done in this work, one would necessarily encounter the strong coupling of the transverse modes. However, by correcting the decomposition non-linearly, one can obtain that only the longitudinal modes become strongly coupled.
Let us apply the procedure found in [39] to cure the strong coupling of the transverse modes to this case as well. We can notice that at the strong coupling scale $L_{\beta str}$, both terms are of the same order in the equation of motion for the tensor modes:
$\Box h_{ij}^{T}\sim-\frac{\beta}{M_{pl}^{2}}P_{ijkl}^{T}\Box\left(\chi_{,k}\chi_{,l}\right),$ (53)
However, the form of this equation suggests that the tensor modes should be transformed:
$h_{ij}^{T}=\bar{h}_{ij}^{T}-\frac{\beta}{M_{pl}^{2}}P_{ijkl}^{T}\left(\chi_{,k}\chi_{,l}\right).$ (54)
This form of the introduction of new couplings suggests that writing the metric perturbations in a manifestly covariant way, also known as the disformal transformation, as
$g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}-\frac{\beta}{M_{pl}^{2}}A_{\mu}A_{\nu}$ (55)
could remove the problematic coupling between the tensor and longitudinal modes. In this section, we will explore if this will indeed be the case. In particular, the goal of the following analysis is to show that with the above-modified decomposition, both tensor and transverse modes will be weakly coupled at all scales up to the Planck length, while the longitudinal modes will become strongly coupled once the scale $L_{\beta str}$ is reached.
In order to show this, let us study the longitudinal modes in quartic order. With (55), we find the following Lagrangian density:
$\mathcal{L}=\mathcal{L}_{h}+\mathcal{L}_{A}+\mathcal{L}_{Ah}$ (56)
$\begin{split}\mathcal{L}_{h}=&\frac{M_{pl}^{2}}{4}\left(2h^{\mu\nu}_{,\mu}h_{\alpha\nu}^{,\alpha}-h_{\mu\nu,\alpha}h^{\mu\nu,\alpha}-2h^{\mu\nu}_{,\mu}h_{,\nu}+h_{,\mu}h^{,\mu}\right)\\ &+\frac{M_{pl}^{2}}{2}\left(-\frac{1}{4}h^{\alpha\beta}h_{\mu\nu}h^{\mu\nu}_{,\alpha\beta}-\frac{1}{8}h_{\mu\nu}h^{\mu\nu}h^{\alpha\beta}_{,\alpha\beta}-\frac{1}{4}h^{\alpha\beta}h_{\beta\nu}h_{\alpha,\delta}^{\nu,\delta}-h^{\alpha\beta}h_{\alpha,\mu}^{\mu}h^{\nu}_{\beta,\nu}\right.\\ &\left.+\frac{1}{2}h^{\alpha\beta}h_{\beta\nu,\mu}h^{\mu,\nu}_{\alpha}-h^{\alpha\beta}_{,\mu}h_{\alpha}^{\mu}h^{\nu}_{\beta,\nu}+\frac{1}{8}hh_{\mu\nu}h^{\mu\nu,\gamma}_{,\gamma}+\frac{3}{16}h_{\mu\nu}h^{\mu\nu}h_{,\gamma}^{,\gamma}+h^{\alpha\beta}h_{\alpha,\mu}^{\mu}h_{,\beta}\right.\\ &\left.+\frac{1}{2}hh^{\mu\nu}_{,\mu}h^{\alpha}_{\nu,\alpha}-\frac{1}{4}hh_{\alpha\nu,\mu}h^{\alpha\mu,\nu}+\frac{1}{4}hh^{\alpha\beta}h_{,\alpha\beta}-\frac{1}{16}h^{2}h^{,\mu}_{,\mu}\right),\end{split}$
$\begin{split}\mathcal{L}_{A}=&-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{m^{2}}{2}A_{\mu}A^{\mu}\\ &+\frac{\beta^{2}}{8M_{pl}^{2}}\left[\partial_{\mu}\left(A_{\alpha}A_{\beta}\right)\partial^{\mu}\left(A^{\alpha}A^{\beta}\right)+\partial_{\mu}\left(A_{\alpha}A^{\alpha}\right)\partial^{\mu}\left(A_{\beta}A^{\beta}\right)-2\partial_{\mu}\left(A^{\mu}A^{\nu}\right)\partial_{\alpha}\left(A^{\alpha}A_{\nu}\right)\right]\end{split}$
and
$\begin{split}\mathcal{L}_{hA}=&-\frac{\beta}{4}A_{\gamma}A^{\gamma}\left(h^{\mu\nu}_{,\mu\nu}-\Box h\right)\\ &+\frac{1}{2}h^{\mu\alpha}F_{\mu\nu}F_{\alpha}^{\;\nu}+\frac{m^{2}}{2}h^{\mu\nu}A_{\mu}A_{\nu}+\frac{1}{2}h\left(-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{m^{2}}{2}A_{\mu}A^{\mu}\right)\end{split}$
Let us now study the above Lagrangian density in terms of components. Our goal is to find the leading order contributions for all of the modes. As we will see, the cubic coupling between $\phi$, $\psi$, and $S_{i}$ that appears in $\mathcal{L}_{h}$ will introduce quartic order terms that are only mixed between the vector modes and the tensor ones. Since we are only interested in the longitudinal modes at quartic order, we can (safely) neglect these terms.
Moreover, the couplings between $\phi$, $\psi$, and $S_{i}$ and the vector modes in the last line of $\mathcal{L}_{hA}$ can be neglected as well. One can show that the longitudinal mode in these relations enters in combination with the $A_{0}$ component, such that once the temporal constraint is resolved, it is multiplied by the mass squared, making such terms sub-dominant.
Therefore, the most important terms are given by:
$\begin{split}\mathcal{L}\sim&M_{pl}^{2}\left(2\phi\Delta\psi-3\dot{\psi}\dot{\psi}-\psi\Delta\psi\right)-\frac{M_{pl}^{2}}{4}S_{i}\Delta S_{i}-\frac{M_{pl}^{2}}{8}\partial_{\mu}h_{ij}^{T}\partial^{\mu}h_{ij}^{T}\\ &-\frac{M_{pl}^{2}}{4}\left(\frac{1}{2}h_{ij}^{T}h_{ik}^{T}h_{jk,\mu}^{T,\mu}+\frac{1}{2}h_{ij}^{T}h_{kl}^{T}h_{kl,ij}^{T}-h_{ij}^{T}h_{il,k}^{T}h_{jk,l}^{T}\right)\\ &+\frac{1}{2}\left[A_{0}\left(-\Delta+m^{2}\right)A_{0}+2A_{0}\Delta\dot{\chi}-\left(\dot{\chi}\Delta\dot{\chi}-m^{2}\chi\Delta\chi\right)\right.\\ &\left.+\left(\dot{A_{i}^{T}}\dot{A_{i}^{T}}-A_{i,j}^{T}A_{i,j}^{T}-m^{2}A_{i}^{T}A_{i}^{T}\right)\right]\\ &+\frac{\beta}{4}\left(A_{0}^{2}-A_{i}A_{i}\right)\left(2\Delta\phi+6\ddot{\psi}4-\Delta\psi\right)+\frac{1}{2}h_{ij}^{T}\left[\left(A_{k,i}^{T}-A_{i,k}^{T}\right)\left(A_{k,j}^{T}-A_{j,k}^{T}\right)-\dot{A}^{T}_{i}\dot{A}^{T}_{j}\right]\\ &+\frac{\beta^{2}}{8M_{pl}^{2}}\left[\partial_{\mu}\left(A_{\alpha}A_{\beta}\right)\partial^{\mu}\left(A^{\alpha}A^{\beta}\right)+\partial_{\mu}\left(A_{\alpha}A^{\alpha}\right)\partial^{\mu}\left(A_{\beta}A^{\beta}\right)-2\partial_{\mu}\left(A^{\mu}A^{\nu}\right)\partial_{\alpha}\left(A^{\alpha}A_{\nu}\right)\right]\end{split}$ (57)
Let us now study the equations of motion. To the leading order, by varying with respect to $\phi$ and $\psi$, we find:
$\phi\sim-\psi\sim\frac{\beta}{4M_{pl}^{2}}\left(A_{0}A_{0}-A_{i}A_{i}\right)$ (58)
and
$S_{i}\sim 0.$ (59)
The $A_{0}$ component of the vector field satisfies the following equation:
$\begin{split}(-\Delta+m^{2})A_{0}\sim&-\Delta\dot{\chi}-3\beta A_{0}\Box\psi-\frac{\beta^{2}}{M_{pl}^{2}}\left[-A_{0}\Delta(A_{0}A_{0})+A_{0}\partial_{0}\partial_{i}(A_{0}\chi_{,i})\right.\\ &\left.+\frac{1}{2}\chi_{,i}\partial_{i}\partial_{0}(A_{0}A_{0})-\frac{1}{2}A_{0}\partial_{0}^{2}(\chi_{,i}\chi_{,i})+\frac{1}{2}\chi_{,i}\Delta(A_{0}\chi_{,i})\right.\\ &\left.-\frac{1}{2}\chi_{,i}\partial_{i}\partial_{k}(A_{0}\chi_{,k})+\frac{1}{2}A_{0}\Delta(\chi_{,i}\chi_{,i})-\frac{1}{2}\chi_{,i}\partial_{0}\partial_{j}(\chi_{,i}\chi_{,j})\right]\end{split}$ (60)
By solving this equation perturbatively, and substituting it into the equation of motion for the longitudinal modes, we find that the leading corrections to the longitudinal modes satisfy:
$\begin{split}m^{2}(-\Box+m^{2})\chi^{(2)}\sim&\frac{\beta^{2}}{M_{pl}^{2}}\left[\frac{9}{2}\Delta\chi\dot{\chi}_{,i}\dot{\chi}_{,i}+11\dot{\chi}\dot{\chi}_{,i}\Delta\chi_{,i}-\dot{\chi}\Delta\dot{\chi}\Delta\chi-\dot{\chi}\dot{\chi}_{,ij}\chi_{,ij}+2\dot{\chi}\dot{\chi}\Delta^{2}\chi\right.\\ &\left.+\dot{\chi}\Delta\dot{\chi}_{,i}\chi_{,i}+\dot{\chi}_{,i}\dot{\chi}_{,ij}\chi_{,j}-\frac{7}{4}\Delta\chi\Delta\chi\Delta\chi-\frac{7}{4}\Delta\chi\chi_{,ij}\chi_{,ij}-\frac{3}{2}\Delta\chi\Delta\chi_{,i}\chi_{,i}\right.\\ &\left.-\frac{3}{2}\chi_{,k}\chi_{,ji}\chi_{ijk}-\chi_{,i}\Delta\chi_{,j}\chi_{,ij}-\chi_{,j}\chi_{,ik}\chi_{,jk}-\chi_{,j}\chi_{,k}\Delta\chi_{,jk}\right]\end{split}$ (61)
It can be easily shown that the first-order corrections $\chi^{(1)}$ will in this case be subleading. Similarly to the previous sections, we can estimate the leading order corrections to the longitudinal modes as:
$\chi^{(2)}\sim\frac{\beta^{2}}{M_{pl}^{2}m^{5}L^{7}}$ (62)
These become of the same order as the linear term $\chi^{(0)}$ precisely at the strong coupling scale that we have previously found:
$L_{\beta str}\sim\left(\frac{\beta}{M_{pl}m^{2}}\right)^{1/3}.$ (63)
At this scale, the most dominant term at the level of Lagrangian density for the longitudinal modes is of the form $\frac{\beta^{2}}{M_{pl}^{2}L^{6}}\chi^{4}$. This suggests that beyond the strongly coupled scale, the minimal amplitude of the longitudinal mode is given by:
$\delta\chi\sim\sqrt{\frac{M_{pl}L}{\beta}}$ (64)
In contrast to the longitudinal modes, we can notice that the tensor and transverse modes are weakly coupled. Their leading contribution comes in the cubic order, and is of the form:
$\begin{split}(-\Box+m^{2})A_{l}^{T}=P_{li}^{T}&\left[\frac{\beta}{2}\left(A_{j}^{T}+\chi_{,j}\right)\Box h_{ij}^{T}+\dot{h}_{ij}^{T}\dot{A}_{j}^{T}\right.\\ &\left.-h_{ij}^{T}\left(A_{i,jk}^{T}-A_{j,ik}^{T}\right)+\partial_{k}\left(h_{ij}^{T}\left(A_{k,j}^{T}-A_{j,k}^{T}\right)\right)\right]\end{split}$ (65)
for the transverse modes, and
$M_{pl}^{2}\Box h_{ij}^{T}\sim-2(A_{k,i}^{T}-A_{i,k}^{T})(A_{k,j}^{T}-A_{j,k}^{T})+2\dot{A}_{i}^{T}\dot{A}_{j}^{T}+\mathcal{O}\left((h_{ij}^{T})^{3}\frac{M_{pl}^{2}}{L^{2}}\right)$ (66)
Here, we have omitted the contribution of the longitudinal modes, which is now subdominant. By taking into account that the minimal amplitude of quantum fluctuations for the tensor modes is singular in the Planck mass, it is clear that the two types of modes will become strongly coupled only at the Planck scale, before which the perturbative approach still holds.
5 Resolving the Runaway Modes
Recently, it was reported in [26] that the non-minimally Proca theory in curved space-time has a runaway production mechanism – mechanism, in which at high energies, modes with negative frequency appear. Following the procedure of [26], let us consider again the action, with the background metric given by:
$ds^{2}=a^{2}(\eta)\eta_{\mu\nu}dx^{\mu}dx^{\nu}.$ (67)
$\begin{split}S_{Pnmin}=\int d^{4}x\sqrt{-g}&\left(-\frac{1}{4}g^{\mu\alpha}g^{\nu\beta}F_{\mu\nu}F_{\alpha\beta}-\frac{m^{2}}{2}g^{\mu\nu}A_{\mu}A_{\nu}\right.\\ &\left.-\frac{1}{2}\alpha Rg^{\mu\nu}A_{\mu}A_{\nu}-\frac{1}{2}\beta R^{\mu\nu}A_{\mu}A_{\nu}\right).\end{split}$ (68)
The above action can be written as:
[…]
where
$M_{T}^{2}=a^{2}\left[m^{2}+\alpha R-\beta\left(3H^{2}-\frac{1}{2}R\right)\right],$ (70)
and
$M_{S}^{2}=a^{2}\left[m^{2}+\alpha R+\beta\left(H^{2}+\frac{1}{6}R\right)\right].$ (71)
Here, $F_{0i}=A_{i}^{\prime}-A_{0,i}$, where the prime denotes the derivative with respect to the conformal time, and $H=\frac{a^{\prime}}{a^{2}}$ is the Hubble parameter. As before, the $A_{0}$ component is not propagating. By varying the action with respect to it, we find the following constraint:
$(-\Delta+M_{T}^{2})A_{0}=-\Delta\chi^{\prime}.$ (72)
By substituting its solution back to the action, and going to the momentum space for the longitudinal modes:
$\chi(\vec{x},\eta)=\int\frac{d^{3}k}{(2\pi)^{3/2}}e^{i\vec{k}\vec{x}}\chi_{k}(\eta)$ (73)
we find the following action for the longitudinal modes:
[…]
By canonically normalizing the field according to:
$\chi_{nk}=\sqrt{\frac{M_{T}^{2}k^{2}}{k^{2}+M_{T}^{2}}}\chi_{k},$ (75)
we find, in the limit when $k^{2}\gg M_{T}^{2},M_{S}^{2}$:
[…]
As pointed out in [26], if we assume that $M_{T}^{2}>0$, in order to avoid the negative kinetic term for the longitudinal modes, then, the case $M_{S}^{2}<0$ leads to the modes with negative frequency. These are called the runaway modes. In other words, with such a choice, one is able to produce modes with arbitrarily large momentum.
One should note though, that for this choice, the transverse modes have negative mass. Their Lagrangian density is given by:
$\mathcal{L}_{T}=\frac{1}{2}\left(A_{i}^{T^{\prime}}A_{i}^{T^{\prime}}-A_{i,j}^{T}A_{i,j}^{T}-M_{S}^{2}A_{i}^{T}A_{i}^{T}\right).$ (77)
We can notice that the origin of the runaway case is the $\beta$ coupling, due to which $M_{T}^{2}\neq M_{S}^{2}$. However, this is precisely the interaction between the Ricci tensor and the vector modes, that was yielding unphysical results in the flat case. Following the previous procedure which resolves the strong coupling of the tensor modes, it is thus natural to introduce the disformal coupling to the metric with the transformation that we have performed before:
$g_{\mu\nu}=\tilde{g}_{\mu\nu}-\frac{\beta}{M_{pl}^{2}}A_{\mu}A_{\nu},\qquad\text{where}\qquad\tilde{g}_{\mu\nu}=a^{2}\eta_{\mu\nu}$ (78)
is the new metric. By performing this substitution, we find that the action (1) becomes:
$\begin{split}S=\int d^{4}x\sqrt{-\tilde{g}}&\left[\frac{M_{pl}^{2}}{2}\tilde{R}-\frac{1}{4}\tilde{g}^{\mu\alpha}\tilde{g}^{\nu\beta}F_{\mu\nu}F_{\alpha\beta}-\frac{m^{2}}{2}\tilde{g}^{\mu\nu}A_{\mu}A_{\nu}\right.\\ &\left.-\frac{1}{2}(\alpha+\beta)\tilde{R}\tilde{g}^{\mu\nu}A_{\mu}A_{\nu}+\mathcal{O}\left(\frac{\beta^{2}}{M_{pl}^{2}}A^{4}\tilde{R}\right)\right]\end{split}$ (79)
in the lowest order, while the last term represent higher order terms that involve couplings between the vector fields with the Ricci scalar or Ricci tensor. This guarantees that now $M_{S}^{2}=M_{T}^{2}$, and thus the runaway case disappears.
6 Conclusion
In this work, we have studied Proca theory, non-minimally coupled to gravity. The cosmological aspects of this theory have recently drawn a lot of attention, making it an intriguing possibility to underline the gravitational production of dark matter and a possible candidate for generating primordial magnetic fields and inflationary candidates. By studying this theory in flat space-time, with metric perturbations also taken into account, we have found a somewhat surprising result. On the one hand, the interaction of the vector field with the Ricci scalar results in the strong coupling of the longitudinal modes. This is relatively natural to expect, as one obtains self-interaction of the longitudinal modes at the quartic order. Beyond it, the vector and tensor modes remain weakly coupled, up to the Planck length, due to the amplitude of quantum fluctuations for the tensor modes. On the other hand, the coupling of the vector modes with the Ricci tensor indicates that both tensor and longitudinal modes become strongly coupled at the same Vainsthein scale. This, however, is an unphysical result – it would imply that if one allows for the non-minimal coupling in nature, the gravitational waves would quickly lose their linear propagator. Moreover, a recent study has found a new class of the negative frequency modes in curved space-time, known as the runaway modes [26]. In this work, we have shown that its origin lies precisely in the coupling, which is problematic even in the flat background.
Nevertheless, by carefully analyzing the theory at this strong coupling scale, we have found that the tensor modes have to be transformed. A similar scenario was found also in massive Yang-Mills theory, with mass added by hand, where with standard decomposition, the transverse modes were becoming strongly coupled. However, by correcting their definition, it was shown that ultimately, these modes remain weakly coupled at all scales.
In the case of non-minimally coupled Proca theory, we have found that the introduction of corresponding disformal coupling in the theory simultaneously removes the inconsistent interaction by introducing the necessary coupling. This leads us to the self-interaction of the longitudinal modes at the quartic order, becoming a necessary step to make the theory safe. Notably, the same transformation removes the problematic runaway modes, while at the same time generating further self-interactions of the vector field. So it will be interesting to study how the predictions made for the original theory change once this is taken into account, if the runaway could re-appear once the theory is quantized, and to explore further aspects of the strong coupling in curved space-time for this type of theory. In particular, it would be intriguing to see if such a possibility could resolve the well-known ghost, that appears for $M_{T}^{2}<0$.
One potential curiosity might be to investigate the theory in the presence of external matter. For this, one has two possibilities – either to first couple the matter with the old metric, or couple it only after the disformal transformation has been performed. Intuitively, one can notice that even in the kinetic term the scalar will generate quartic interactions in this case. In contrast, if the external matter is added aposteriori, the possibility of the scalar field feeling the interaction could vanish more safely. While it would be interesting to check this analysis for both scenarios, a realistic perspective is that the matter should be minimally coupled to the transformed metric, as in this frame tensor modes are weakly coupled, which is in an agreement with what we observe today. We will leave further aspects of this investigation for the upcoming works. Finally, one should note that while in Proca theory, the longitudinal modes are the source of the singularities in mass within the perturbative regime, in the Kalb-Ramond field, this role is taken by the vector modes [50]. Therefore, it will be interesting to also extend this analysis to this case, and study its consequences.
Acknowledgements
I am grateful to Misao Sasaki for illuminating discussions and comments on the draft of the paper. I would also like to thank Edward "Rocky" Kolb for communicating his work on the Proca theory, and for very useful discussions. In addition, I would like to thank Elisa G. M. Ferreira, Shinji Mukohyama, Ryo Namba, and Ippei Obata for very useful discussions. The work of A. H. was supported by the World Premier International Research Center Initiative (WPI), MEXT, Japan.